Solomon Entertainment sells souvenir T-shirts at each rock concert that it sponsors. The shirts cost $6 each. Any excess shirts can be returned to the manufacturer for a full refund of the purchase price. The sales price is $26 per shirt Required a. What are the total cost of shirts and cost per shirt if sales amount to 6,000, 6,500, 7000, 7500, or 8,000? Answer is complete but not entirely correct. 7,000 $ 36,000 39,000 42,000 45,00048,000 6,000 6,500 7,500 8,000 Total cost of shirts Cost per shirt 36,000 39,00042,000 45,000 48,000
Answer #1

Solution:- In order to calculate the total cost of good, numbers of units multipled by Cost per unit.

Total cost=(number of units produce×Cost per unit)

     And In order to calculate Cost per unit, Total cost divided by number of units.

Cost per unit=(Total cost of good÷numbers of good produce at that cost)

No.of shirts Sold Total cost of shirts Cost per shirt
6,000 $(6,000×6)=$36,000 $6
6,500 $(6,500×6)=$39,000 $6
7,000 $(7,000×6)=$42,000 $6
7,500 $(7,500×6)=$45,000 $6
8,000 $(8,000×6)=$48,000 $6

Cost per shirt is given $6.

On the other hand if Calculating the Cost per shirt is:-

Cost Per shirt=(Total cost of Shirt÷no. Of shirts)

=$(36,000÷6,000)

=$6